### v1

# In-Context Learning on a Budget: A Case Study in Named Entity Recognition

Abstract
Few shot in-context learning (ICL) typically assumes access to large annotated training sets. However, in many real world scenarios, such as domain adaptation, there is only a limited budget to annotate a small number of samples, with the goal of maximizing downstream performance. We study various methods for selecting samples to annotate within a predefined budget, specifically focusing on the named entity recognition (NER) task, which has real-world applications, is expensive to annotate, and is relatively less studied in ICL setups. Across different models and datasets, we find that a relatively small pool of annotated samples can achieve results comparable to using the entire training set. Moreover, we discover that random selection of samples for annotation yields surprisingly good performance. Finally, we observe that a diverse annotation pool is correlated with improved performance. We hope that future work adopts our realistic paradigm which takes annotation budget into account.
In-Context Learning on a Budget:
A Case Study in Named Entity Recognition
Uri Berger1,2, Tal Baumel3, Gabriel Stanovsky1 1School of Computer Science and Engineering, The Hebrew University of Jerusalem 2School of Computing and Information Systems, University of Melbourne 3Microsoft Corporation {uri.berger2, gabriel.stanovsky}@mail.huji.ac.il tal.baumel@microsoft.com
1 Introduction
In-context learning (ICL) has emerged as a highly efficient and robust method for various textual tasks. In this paradigm, a large language model (LLM) is exposed to a small number of annotated samples, termed demonstration examples, which are provided as part of the prompt, before the sample which the model is required to annotate, which we will refer to as inference sample henceforth. While the reasons for ICL’s success are still contested Min et al. (2022); Liu et al. (2022), it has been observed that ICL prompts commonly outperform zero-shot prompts, where no demonstration examples are provided Brown et al. (2020).
Furthermore, a recent line of work has found that the choice of demonstration examples can lead to improved results over random demonstration selection. For example, Liu et al. (2022) found that choosing the nearest neighbors of the inference sample in the training set leads to improvements over random demonstration selection on 6 tasks, such as sentiment analysis or question answering. In all of these, the demonstration examples are chosen from large annotated training sets, ranging from 3.5K samples up to 78K samples.
In this work, we address the following research question: How can we maximize ICL performance on a given annotation budget? This question is particularly relevant for real-world domain adaptation settings, where a large pool of annotated samples is unavailable for selecting demonstration examples. Instead, there are large sets of unannotated samples (e.g., raw text in the target domain), and a fixed budget to annotate only a small portion of them. As depicted in Figure 1, we define the task as pool selection, i.e., selecting a small pool of $k$ examples out of a large corpus of raw texts. These samples are annotated and serve as the available pool for demonstration examples.
Figure 1: Our proposed approach for ICL on a budget, illustrated in four steps: (1) we assume a large pool of raw train and potentially test texts in the target domain; (2) a pool selection strategy chooses a subset of $k$ train texts to maximize downstream ICL performance; (3) annotations are collected for the selected pool; (4) an inference prompt is instantiated by choosing the nearest examples in the pool to the inference sample. We focus on step (2), experimenting with various pool selection strategies.
We implement various methods for pool selection, e.g., clustering the train set and selecting a representative example from each cluster. Within the scope of this paper, we focus on the task of named entity recognition (NER), as it has various applications in real-world downstream tasks (e.g., finance, biomedical (Pakhale, 2023)), it is relatively understudied in the context of ICL Chen et al. (2023), and is expensive to annotate due to the need for linguistic expertise and domain-specific knowledge Chen et al. (2015).
We evaluate several state-of-the-art LLMs on two NER benchmarks. We first replicate previous findings of the strong effect of choice of demonstrations on performance. Second, we find that a relatively small pool ($\sim$200 samples) allows LLMs to perform over 88% as when demonstrations are selected from the full training set. However, when compared to NER state-of-the-art models, the performance gap remains considerable in some cases. Additionally, we find that in some scenarios randomly selecting samples for the annotation pool performs similarly to more carefully designed methods. Finally, we show that across models and datasets, a more diverse pool is correlated with improved performance.
We hope that our paradigm is adopted in future work in order to report more realistic ICL performance, and explore new methods for sample selection for other tasks and domains.
2 ICL on a Budget
Here we propose a conceptual framework for ICL in a realistic domain adaptation setting (depicted in Figure 1), where there are no apriori annotated datasets for the target task. Instead, we assume that there is a large corpus of raw texts in the target domain, and a fixed budget for annotating a small portion of them, such that they can serve as potential demonstration exemplars during inference. Intuitively, the goal of the annotation process is to maximize downstream ICL performance.
Below we formalize the task of pool selection, and describe 5 selection strategies, which aim to maximize different aspects, e.g., coverage of the training set versus coverage of the test set. In the following section we evaluate these approaches for the NER task.
2.1 Pool Selection: Task Definition
Formally, a pool selection strategy is a function:
$S:\big{(}\mathcal{P}(\mathcal{D})\cup\mathcal{P}(\mathcal{T})\big{)}\mapsto\mathcal{D}^{k}$ (1)
Where $\mathcal{D}$ represents an unannotated “training” distribution (e.g., all texts in a certain domain), $\mathcal{T}$ similarly represents a “test” distribution, $\mathcal{P}(\mathcal{D})$ and $\mathcal{P}(\mathcal{T})$ are the power sets of $\mathcal{D}$, $\mathcal{T}$ respectively, and $k>0\in\mathbb{N}$ is the annotation budget, i.e., the number of samples to annotate. Intuitively, $S$ maps raw train and test sets to $k$ train samples (the pool), which are then annotated. In all that follows we denote the input train set by $D\in\mathcal{P}(\mathcal{D})$ and test set by $T\in\mathcal{P}(\mathcal{T})$. Typically, $k<<|D|$ and $k<<|T|$, indicating that the annotation budget for a new domain can only annotate a small portion of its available texts.
Furthermore, we assume a similarity function $\phi$:
$\phi:\big{(}(\mathcal{D}\cup\mathcal{T})\times(\mathcal{D}\cup\mathcal{T})\big{)}\mapsto\mathbb{R}$ (2)
In the scope of this work (similar to previous work (Liu et al., 2022)) samples are embedded into $\mathbb{R}^{m}$, where $m$ is the text embedding dimension, and $\phi$ is the cosine similarity of these vectors. In particular, we use a sentence transformer (Reimers and Gurevych, 2019) trained over MPNet (Song et al., 2020).
2.2 Pool Selection Strategies
Below we describe 5 pool selection strategies which follow the definition in Equation 1. Each of these may be better suited for different settings and downstream applications.
The first two strategies (largest mean and cluster test) assume that the test set distribution $\mathcal{T}$ is known. These may be best for a well-delineated downstream application. In contrast, the last three strategies (cluster train, vote-K, and random) do not assume the distribution on which the model will be tested (i.e., $\mathcal{T}$ is unknown) and hence apply more widely.
Largest mean.
For each sample $d\in D$, compute its mean similarity to all samples in $T$:
[…]
Next, choose the $k$ samples from $D$ with the largest mean_sim value. Intuitively, this approach seeks the $k$ train samples which are closest on average to the test set.
Cluster test.
Cluster $T$ into $k$ clusters using $\phi$ as the similarity method. For each cluster center choose the most similar sample from $D$. Intuitively, this approach tries to maximize the coverage of the test set distribution.
Cluster train.
Cluster $D$ into $k$ clusters and for each cluster center choose the most similar sample from $D$. In contrast to the cluster test strategy, cluster train aims to maximize the coverage of the expected training distribution. This strategy was proposed in Chang et al. (2021) for selecting examples for fine-tuning.
Vote-k (Su et al., 2022).
Selects a $k$-sized subset of D such that the samples are diverse and dissimilar from one another, through a two-step process which uses the LLM’s confidence to bucket the different samples. See Su et al. (2022) for more details.
Random.
Randomly select $k$ samples from $D$.
3 Evaluation
Figure 2: F1 score for strict matching using Claude 3 Haiku (left), GPT4 (middle) and Gemini 1.5 Pro (right) on Ontonotes (top) and UNER (bottom), as a function of the size of sample pool. In methods with a random component we run 3 trials and plot error bars showing the standard deviation. Oracle: using the full training set as the pool.
We focus on the NER task, as it is both applicable in real-world settings, while traditionally expensive to annotate at scale. Importantly, as is commonly done in active learning scenarios (e.g., Shen et al., 2018; Liu et al., 2020), we use existing NER datasets in various domains, allowing us to simulate the process shown in Figure 1, where instead of collecting manual annotations, we sample from the existing annotations.
3.1 Experimental Setup
Datasets and models.
We use two English datasets for NER: Ontonotes 5.0 Hovy et al. (2006),111 We use the HuggingFace version with the english_v12 configuration. and the recently introduced Universal NER (UNER; Mayhew et al., 2023). Both datasets contain texts from various domains, including news, conversational, weblogs, web forums, and more. Due to budget constrains we limit the size of each test set by randomly sampling 1000 samples.
We experiment with a diverse set of 9 different LLMs, ranging from open to closed models in various parameter sizes: Llama-2 Touvron et al. (2023), Mistral Jiang et al. (2023), Starling Zhu et al. (2023), Vicuna Chiang et al. (2023), Mixtral Jiang et al. (2024), phi-2 Javaheripi et al. (2023), GPT-4 Achiam et al. (2023), Claude 3 Haiku,222www.anthropic.com/news/claude-3-family and Gemini 1.5 Pro Reid et al. (2024).
Prompt and evaluation metric.
We prompt the models by describing the NER task and the types of entities, followed by 5 demonstration examples and the current inference sample. Following Zhao et al. (2021), we sort the examples such that the most similar example is the last one, based on the similarity function $\phi$ described in Equation 2. See Appendix A for the full prompt template.
We evaluate all models with the standard strict matching evaluation metric for NER Segura-Bedmar et al. (2013), where a predicted entity is considered correct if it matches both exact span boundaries and entity type.
Pool sizes.
We experiment with sample pools of size 0.1% - 10% of the maximum pool size, defined as the number of unique samples used when considering the full training set as the sample pool.
Oracle.
To better understand the performance of different pool selection strategies, we report the performance of using the full training set for each dataset and model, compared to our budget-constrained approach.
3.2 Results
Results for the three best-performing models (Claude 3 Haiku, GPT-4, and Gemini 1.5 Pro) on both of our datasets are presented in Figure 2. The other smaller models we test are not able to produce outputs of the requested format in more than 50% of the cases, and hence cannot be meaningfully compared against these models, which adhere to the correct format in roughly 97% of the cases. [^3] Gemini 1.5 Pro was unable to produce the requested format for small pool sizes in UNER, and we omit these results. We now discuss findings reflected in these results and conduct further analysis.
The choice of few-shot examples matter in NER.
We observe a large variation in performance when selecting different demonstration examples in all configurations. While this was observed in other tasks (Zhang et al., 2022), to the best of our knowledge, this is the first time this was shown for NER.
Most pool selection methods perform similarly, random is surprisingly good.
None of the five methods demonstrated a significant improvement or decline in performance compared to the others, with the exception of largest mean, which showed notably lower scores. Surprisingly, we note that random performs similarly to other methods. We note that this is true even for selection strategies that are allowed access to the test set distribution.
Very small pool sizes can approximate the full training corpus.
With the exception of the low-performing largest mean, all sampling methods require a pool of only 220 samples for Ontonotes, and 147 samples for UNER, to achieve over 88% of the oracle performance across all configurations.
ICL on a budget approaches state-of-the-art on UNER, but lags behind on Ontonotes.
We compare the ICL results to state-of-the-art results in Table 1 in Appendix B. In the well studied Ontonotes, fine-tuned methods vastly outperforms using a limited annotation budget. In the relatively new UNER dataset, using 147 samples achieves more than 90% of the state-of-the-art.
The label diversity in a sample pool is correlated with its ICL performance.
Min et al. (2022) find that the coverage of the label space by the demonstration examples has a strong effect on performance. Drawing inspiration from their findings, we examine whether the diversity in the labels of demonstration examples is correlated with performance. To this end, for each dataset, pool selection method, and pool size, we count how many instances of each entity were presented in the demonstration examples in the sample pool, and compute the entropy of these counts as a proxy for diversity. For each model, we then compute the Pearson correlation between these entropy values and the model’s scores. We find that the diversity of labels in the demonstration examples is indeed significantly correlated with performance across all models and datasets. See Table 2 in Appendix C for the correlation values.
4 Conclusion
We proposed the framework of ICL on a budget and studied different methods for pool selection, focusing on NER. We hope this work will inspire more realistic assumptions on the amount of labeled data used in different ICL settings. In future work we will extend this analysis to additional tasks with high annotation costs.
5 Limitations
We tested our ICL on a budget approach on a single task (NER), because it has real-world applications, high annotation cost, and it was relatively less studied in the context of ICL. It is possible that other tasks will show different trends, hence we stress that the contribution here is the methodological approach, rather than advocating for one particular sampling strategy.
Appendix A Prompt Template
In this section we describe the prompt template we use in our experiments.
We first describe the task:
You are an NER classifier that identifies the following entities: Person __PER__, Organization __ORG__, Geo-Political __GPE__, Location __LOC__, Facility __FAC__, Work-of-Art __WOA__, Event __EVE__, Product __DUC__, Language __ANG__ use angle brackets to tag in-line, please don’t include any additional information other than the annotated sentence and keep original spacing.
Next, we add the demonstration examples. For GPT and Claude we add the demonstration examples as follows. For each demonstration example, we use LangChain’s444langchain.com HumanMessage class for the original sentence, followed by an AIMessage for the annotated sentence. Finally, we add the inference sample as a HumanMessage.
In Gemini, for which these classes are not implemented, we choose a different strategy: for each demonstration example, we add the original sentence followed the separator token [TAGS], and then the annotated sentence. Finally, we add the inference sample, followed by the separator token. Figure 3 demonstrates a prompt for Gemini.
Figure 3: An example prompt used in our study.
Appendix B Comparison to State-of-the-Art
Table  1 presents the percentage in performance of the best performing model, GPT-4 out of the state-of-the-art, in each dataset: Ontonotes (92 strict f1-score;  Li et al., 2020) and UNER (86 strict f1-score; Mayhew et al., 2023).
Ontonotes UNER
Cluster train 75 91
Cluster test 76 93
Largest mean 67 88
Vote-k 78 91
Random 78 91
Table 1: GPT-4 percentage in performance out of the state-of-the-art, when using 5% of the samples used in the fully labeled train set method.
Appendix C Performance-Diversity Correlation
In Section 3.2 we computed the Pearson correlation of performance with diversity, as measured by the entropy of counts of entities. Table 2 presents the correlations.
GPT Claude Gemini
Ontonotes 0.54 0.36 0.46
UNER 0.36 0.40 0.44
Table 2: Pearson correlation of performance with label diversity in the sample pool, as measured by the entropy of entities. All results are significant.